# DrivAer Transformer: A high-precision and fast prediction method for vehicle aerodynamic drag coefficient based on the DrivAerNet++ dataset

### v3

@@ V.2 Discussion
    (a) Basic Model
    (b) Enhanced Model
    Figure 15: Error Distribution Across Different Vehicle Types
    Table 3: Error Rate Comparison between Basic and Enhanced Models.
@@ VI Conclusion
In this paper, we obtained data on the aerodynamic performance of automobiles through numerical simulation and high-fidelity CFD simulation, and constructed a large-scale multimodal dataset, DrivAerNet++, containing different design parameters and aerodynamic coefficients. Eventually, this study established a prediction model DAT based on geometric deep learning and verified the differences of different models without using test set samples. The algorithm accomplishes innovations in the following two points: 1. Utilizing the negative feedback regulation mechanism in the field of automatic control, the idea of error-correcting feedback structure to comprehensively capture the local features of the point cloud; 2. Utilizing the attention module based on the affinity of the channel to help avoid possible redundancy in feature mapping. The conclusions of the study are as follows:
@@
    Using the DrivAerNet++ dataset and the geometric deep learning model, this study is able to efficiently and accurately predict the drag coefficient of a car. The model’s error is controlled within 4% under both the initial validation set and the supplementary validation set, showing high prediction accuracy and greater generalizability to other models relative to the RegDGCNN and DGCNN architectures. Moreover, the modified algorithm can directly process 3D mesh data, and the model does not require additional image rendering or generation of symbolic distance function (SDF), which simplifies the preprocessing steps and improves the computational efficiency and practicality of the model.
@@
    By extending the validation dataset with more data from SUV, MPV and other models, this study enhances the generalization ability of the DAT model. In these extended datasets, the DAT architecture outperforms the traditional RegDGCNN model, showing greater adaptability and higher prediction accuracy. For different air drag coefficient $C_{d}$ value intervals, the prediction error of the DAT model is reduced. This indicates that the model works better in handling new data and improves in the prediction of aerodynamic performance of complex 3D automobile shapes. The results after validation using wind tunnel experiments to obtain accurate data show that the DAT model not only meets the current automobile manufacturing industry’s requirements for prediction accuracy, but also demonstrates promising applications on large-scale multimodal datasets.
@@
    Although the DAT model demonstrated excellent performance in experiments, the study also identified some potential problems in the training and optimization process, such as limited computational resources, high data processing complexity, and high cost of wind tunnel testing. Future research should focus on streamlining the data management and loading strategies, improving the efficiency of computational resource usage, and exploring more cost-effective experimental validation methods to promote the popularization of the DAT architecture in practical engineering applications. Through continuous optimization and extension, the model is expected to gain wide and far-reaching influence in automotive aerodynamic performance prediction and other related fields.

### v5

@@ V.5 Discussion
    Table 5: Error Comparison between Basic and Enhanced Models.
@@ VI Conclusion
In this paper, we obtain automotive aerodynamic performance data through high-fidelity CFD numerical simulation, construct a large-scale multimodal dataset DrivAerNet++ containing multiple design parameters and aerodynamic coefficients, and establish a DAT prediction model based on geometric deep learning, and validate the performance differences of different models without using the samples of the test set. The algorithm is innovative in the following two aspects: 1. Utilizing the negative feedback regulation mechanism in the field of automatic control, the idea of error-correcting feedback structure to comprehensively capture the local features of the point cloud; 2. Utilizing the attention module based on the affinity of the channel to help avoid possible redundancy in feature mapping. The conclusions of the study are as follows:
@@
    Utilizing the DrivAerNet++ dataset with the geometric deep learning model, the vehicle drag coefficients can be predicted efficiently, and the model’s errors on both the initial and extended validation sets are controlled within 4%. Compared with RegDGCNN and DGCNN, the DAT model has higher accuracy and generalization ability, and can directly process 3D mesh data without additional image rendering or SDF construction, which improves the efficiency and practicability of pre-processing.
@@
    The generalization ability of the model is enhanced by expanding the validation set by introducing SUV and MPV models. Compared with the RegDGCNN model, the DAT model has lower prediction error in each drag coefficient $C_{d}$ range, and shows better adaptability and complex shape prediction capability. Combined with wind tunnel measurements, DAT not only meets the current industrial accuracy requirements, but also shows the potential to be applied to large-scale multimodal data.
@@
    Despite the excellent performance of DAT, the study still finds that the training and optimization process has the problems of computational resource constraints, data processing complexity, and high cost of wind tunnel tests. Future research should focus on simplifying data management and loading strategies, improving computational efficiency, and exploring more cost-effective validation methods to promote the application of DAT in engineering practice. Through continuous optimization, DAT is expected to have a wide and far-reaching impact on the prediction of automotive aerodynamic performance and related fields.
Supplementary Material
This supplementary document provides an additional visualization supporting the main manuscript. The figure shows the configuration-specific predicted $C_{d}$ values across various subsets of the DrivAerNet++ test set.